A simple pendulum undergoes SHM with time period T.How much time will it take to complete 3/8 oscillations?

Grade:11

1 Answers

Arun
5 years ago
 
Time taken to complete one fourth oscillation T/ 4
so time to complete 1/8th oscillation from to extreme position is obtained from 
y= a/2 = a cos cot = a cos 2 pai / T × t
0r cos Pai /3 = cos 2pai t/ T
t = (T/6) s 
hence time to complete 3/8th oscillation 
= T/4 + T/6 
= 5T /12 
 
 
Regards
Arun